OpenStudy (anonymous):

Write a system of equations with the solution (4,-3)

OpenStudy (jdoe0001):

(4,-3) <---- what does that mean?

OpenStudy (anonymous):

Thats the coordinates I believe

OpenStudy (jdoe0001):

coordinates... for?

OpenStudy (anonymous):

graph

OpenStudy (jdoe0001):

have you done system of equations yet?

OpenStudy (anonymous):

thats what were learning but i have trouble learning that

OpenStudy (jdoe0001):

can you tell where the 2 lines meet?

OpenStudy (anonymous):

(4,-3)

OpenStudy (jdoe0001):

well... what does that mean? well it means that, if we use that for "x" and "y" on BOTH equations, the result will be as indicated by the = sign... so let's check that \(\large \begin{array}{llll} 2x + 3y = -1\implies 2({\color{red}{ 4}})+3({\color{blue}{ -3}})=-1\implies &8-9=-1\\ 5x-8y = 44\implies 5({\color{red}{ 4}})-8({\color{blue}{ -3}})=44\implies &20+24=44 \end{array}\) so it checks out

OpenStudy (jdoe0001):

so to make any other with whatever coordiinates all you do is say .... coordinates (3, 13) so x = 3 y = 13 so I can just plug in whatever in front of it 25x + 3y = 3x-25y = and then plug in the values for "x" and "y" from the coordinates,and get whatever is they're equals to

OpenStudy (anonymous):

Thanks!!

OpenStudy (jdoe0001):

yw